# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ 6.5. Testing and test generation
Tasks related to testing, such as generating test cases, were identified by 28 (5.1%) respondents as having potential legal ramifications. Some respondents identified software testing itself as a shortcoming of generative AI’s capabilities; as $R_{478}$ put it, “[W]ho will test the generated test?” Beyond this, participants asserted that the application of generative AI for testing could lead to legal problems in several ways, including use for penetration testing ($R_{299}$) and security audits ($R_{364}$). The specter of liability also emerged ($R_{314}$): “AI models can automate testing processes, but if automated testing leads to false negatives or positives that result in financial or reputational harm, it may result in legal disputes.” $R_{434}$ also cautioned that “[g]enerating random datasets (csv, image, audio, video, etc.) for software testing may violate someone’s personal rights.”
Finding 30: Developers had concerns related to AI-generated tests, including their completeness and reliability.
Table 7. Main findings for RQ3: Other legal concerns anticipated by developers
6.5 Liability concerns
F25
Developers raised questions about who would be held accountable for bugs and other defects in GenAI
outputs, especially for mission-critical systems such as medical, aerospace, and security. Developers were also
concerned about how would be held responsible for actions that AI agents take autonomously.
6.1 Generating content other than code
F26
The generation of other types of media (images, fonts, videos, audio, etc. and theme/style-sheets can also
pose potential legal challenges for development teams since such assets are often used in product creation.
6.2 Generation of malicious content
F27
Respondents identified the purposeful generation of malicious content as a societal problem and legal
concern going forward.
6.3 Data privacy and information leakage
F28
Developers and other stakeholders did not intend for or anticipate that their data would be collected en masse,
so developers are concerned that private and sensitive information can easily makes its way into training sets.
F29
Generative AI, particularly LLMs, can not only leak information that was contained in the original
training data but can also be used by malicious actors to exploit systems.
6.4 Testing and test generation
F30 Developers had concerns related to AI-generated tests, including their completeness and reliability.
@@ 7.1. Construct validity
Given the design of our study, the results represent developers’ communicated perceptions, which may or may not reflect their or their organizations’ actual practices. We have avoided making any claim connecting such perceptions to actual practices. While we sought to include a wide range of perspectives, the demographic information included in this study was self-reported by respondents and was not independently verified. We also acknowledge that neither the survey respondents nor the interviewees may constitute a representative sample of developers. However, through our coding process, we aimed to report those views and perceptions that were held by several developers.
@@ 7.2. Internal validity
We followed best practices when developing the questionnaire to ensure that the questions were clear and would not bias the participants toward a particular answer. Additionally, we also conducted a small pilot study with graduate students in our research lab to check readability, length, and content. Given that participation in the survey was done on an entirely voluntary basis from a single source of participants, we are aware of the problems introduced by self-selection bias. However, the number of participants gives us confidence that the results capture various developers’ views. In the qualitative analysis of survey responses, we mitigated confirmation bias by having multiple annotators examine the data independently and then reconvene to discuss and resolve any disagreements, rigorously following the best practices of qualitative analysis. Our conclusions were derived from data analysis.
@@ 7.3. External validity
To investigate developers’ perspectives, we rely on GitHub developers who showed interest in repositories of GenAI tools for coding. While our respondents may represent developers working in different domains and organizations, we are also aware that they may overrepresent views typical of open-source ecosystems and underrepresent those of closed-source developers. Although we do not claim generalizability, many respondents work for companies or in education, which may lead to perspectives that vary from those in the open-source community. Moreover, while we had participants from 73 countries spanning 6 continents, nearly a third of respondents came from the U.S., so our results may apply most directly to the U.S. context and legislation.
8. Discussion, Implications, and Conclusions
@@ 8.1. Developing new licenses for software used in AI training
Open-source software (OSS) is used to train GenAI models, and so licensing is an important consideration. Respondents expressed a range of views on the permissibility of using OSS as training data (Section 5.2.1), with some condoning such use (Section 5.2.3) and others believing that OSS developers should take proactive steps if they want their work to be excluded from such use, such as including restrictions in a license (Section 5.2.2). For many years, there has been a standard set of licenses in the OSS community, and drafting new licenses (“license proliferation”) has been discouraged, as it makes compliance tasks more difficult (Bandi, 2019; Wintersgill et al., 2024). However, the text of the most popular OSS licenses was drafted decades ago and did not anticipate technological developments such as GenAI. This means that it may be unclear whether or how the terms of existing OSS licenses apply to scenarios in which OSS projects are used to train, develop, and distribute AI systems.
@@
In the past, updated versions of popular licenses have been drafted to account for new technology (Team, 2021; Stallman, [n. d.]), and AI model-specific licenses (e.g., OpenRAIL (Responsible AI, 2022)) have emerged that impose ethical usage requirements. Thus, the new legal challenges introduced by large-scale training and GenAI may necessitate the drafting of new OSS licenses or the revision of existing licenses so as to provide guidance on these issues. Importantly, since most existing OSS licenses were drafted by developers, who typically do not have legal experience, their terms are sometimes difficult to align with legal doctrine. This has given rise to the less-than-ideal situation in which drafting organizations issue FAQs explaining the intent of their licenses (Wintersgill et al., 2024). If those with legal expertise, preferably with accompanying SE backgrounds, participate in drafting these new licenses, it is possible that ambiguous and unclear language can be avoided to the largest extent possible.
8.2. Disclosing training data and tracking data provenance
Participants expressed a desire that information about the content and provenance of training data be made available for transparency (Sections 5.2.2 and 5.2.3), specifically to fulfill attribution requirements, to allow community checks for security risks or harmful content, to provide a mechanism for creatives to determine if their works have been included, to give users information through which to evaluate tools on ethical and philosophical grounds, and to allow organizations to ascertain legal risk from using tools trained on datasets.
Providing complete information on the content in training material comes with both technical and legal hurdles. It may appear sensible at first to apply proposed DataBOMs (Data Bills of Materials) (Stalnaker et al., 2024; Barclay et al., 2019) to the problem, but the issue becomes one of scale. Unlike more canonical machine learning techniques, LLMs are trained on billions, if not trillions, of data points collected from large swaths of the internet. Such datasets are hard to navigate and even harder to curate (Lee et al., 2023b), and any attempt to provide a complete list of the training data must consider that a work might appear in several different sources. For example, even if a creator’s original work is not directly included in the dataset, it might appear in a StackOverflow post that is included in the dataset. As such, a single DataBOM will likely be insufficient.
Data provenance is similarly complicated. While such functionality has been achieved by newer models, like ChatGPT-4 and Gemini, this relies on Retrieval Augmented Generation (RAG)-style systems (Ma et al., 2024). Achieving correct citations for model generations that do not rely on external data sources accessible through RAG or provided in a prompt is still an open research area.
Since new LLM models can be created by fine-tuning existing base models, it also becomes necessary to appropriately track the provenance and lineage of models. Failure to do so could result in not all training data being known or disclosed. These relationships emerging from the AI supply chain could be tracked using standardized model cards (Mitchell et al., 2019) or AIBOMs (AI Bills of Materials) (Xia et al., 2023; Stalnaker et al., 2024).
Additionally, entities might train or fine-tune a model using proprietary or otherwise sensitive data sources. This creates a tension between transparency and protecting intellectual property or proprietary secrets. Should entities be required to provide traceability and transparency information for the models they release or otherwise make available? Balancing provenance with these data privacy concerns will also be a challenge for future work.
@@ 8.2. Disclosing training data and tracking data provenance
Finally, some respondents called for a right to be removed from training data. Apart from the difficulty of determining whether a data point is included in a training dataset, the further question of what it means to be “removed” and how such removal can be accomplished effectively remain unanswered (Lee et al., 2023a; Yao et al., 2023).
@@ 8.3. New AI advancements will likely amplify legal challenges
Many of the study participants expressed the sentiment that “tools are tools” and, as such, developers should hold the copyright to GenAI output (Section 5.4.2). Much of this understanding is based on the current limitations of existing tools, which, in many cases, require developers to manually correct, edit, or adapt the solution provided by the AI before the code can be used. A few respondents, however, indicated that they anticipated more complicated legal challenges ahead once the technology advances and can produce entire software packages composed of multiple files that largely resemble existing licensed components, a process already under development (Wu, 2024; gpt-pilot, [n. d.]). The question will then be whether developers’ attitudes about generated code ownership will change based on the nature and scope of developers’ own contributions. Likewise, while developers’ intuitions about code usage may currently align, at least to some extent, with U.S. copyright law, there may well be a future misalignment as both the technology and the law evolve. Having a better understanding of developer expectations and perceptions could help guide regulators and decision-makers in this area.
8.4. Concluding remarks
@@
Although there are still many open questions regarding the use of generative AI and copyright, the biggest priority for future research is to address problems related to data provenance. Developers expect tools that can provide them with citations and will not lead them unwittingly into legal trouble. Likewise, they need tools that allow them to document and track code generated by LLMs for transparency and control of legal risk. Given that this paper primarily examines these issues from a perspective grounded in U.S. law, we encourage further research to interpret our results against regulatory efforts elsewhere.

### v3

@@ 6.4. Testing and test generation
Tasks related to testing, such as generating test cases, were identified by 28 (5%) respondents as having potential legal ramifications. Some respondents identified software testing itself as a shortcoming of generative AI’s capabilities; as $R_{478}$ put it, “[W]ho will test the generated test?” Beyond this, participants asserted that the application of generative AI for testing could lead to legal problems in several ways, including use for penetration testing ($R_{299}$) and security audits ($R_{364}$). The specter of liability also emerged ($R_{314}$): “AI models can automate testing processes, but if automated testing leads to false negatives or positives that result in financial or reputational harm, it may result in legal disputes.”
Finding 29: Developers had concerns related to AI-generated tests, including their completeness and reliability.
@@ 7.1. Construct validity
Given the design of our study, the results represent developers’ communicated perceptions, which may or may not reflect their or their organizations’ actual practices. We have avoided making any claim connecting such perceptions to actual practices. While we sought to include a wide range of perspectives, the demographic information included in this study was self-reported by respondents and was not independently verified. Absent a foolproof method of prohibiting duplicate responses, it may have been possible for a participant to submit more than one response. However, no incentive was associated with the survey that would motivate such behavior, and to the best of our knowledge, we received no duplicate responses. We also acknowledge that neither the survey respondents nor the interviewees may constitute a representative sample of developers. To mitigate this threat, through our coding process, we aimed to report those views and perceptions that were held by several developers. Lastly, our respondent sample may overrepresent developers who frequently use GenAI tools and underrepresent those who rarely or never use GenAI tools.
@@ 7.2. Internal validity
We followed best practices when developing the questionnaire to ensure that the questions were straightforward and would not bias the participants toward a particular answer. We also conducted a small pilot study with graduate students in our research lab to check readability, length, and content. Given that participation in the survey was an entirely voluntary basis from a single source of participants, we are aware of the problems introduced by self-selection bias. However, the number of participants gives us confidence that the results capture various developers’ views. In the qualitative analysis of survey responses, we mitigated confirmation bias by having multiple annotators examine the data independently and then reconvene to discuss and resolve any disagreements, rigorously following the best practices of qualitative analysis. Our conclusions were derived from data analysis.
@@ 7.3. External validity
To investigate developers’ perspectives, we rely on GitHub developers who showed interest in repositories of GenAI tools for coding. While our respondents may represent developers working in different domains and organizations, we are also aware that they may overrepresent views typical of open-source ecosystems and underrepresent those of closed-source developers. Although we do not claim generalizability, many respondents work for companies or in education, which may lead to perspectives that vary from those in the open-source community more generally. Moreover, while we had participants from 73 countries spanning six continents, nearly a third of respondents came from the U.S., so our results may apply most directly to the U.S. context and legislation.
8. Discussion and Implications
We distill the key issues highlighted by developers, as reported in our study, and discuss their implications. We also examine potential directions for future research in light of the identified challenges.
@@ 8.1. Developing new licenses for software used in AI training
Because open-source software (OSS) is used to train GenAI models, licensing is an important consideration. Respondents expressed a range of views on the permissibility of using OSS as training data (Finding 5.2.1), with some condoning such use (Finding 5.2.3) and others believing that OSS developers should take proactive steps if they want their work to be excluded from such use, such as including restrictions in a license (Findings 5.2.2 and 5.2.2). For many years, there has been a standard set of licenses in the OSS community, and drafting new licenses (“license proliferation”) has been discouraged, as it makes compliance tasks more difficult (Bandi, 2019; Wintersgill et al., 2024). However, the text of the most popular OSS licenses was drafted decades ago and did not anticipate technological developments such as GenAI. This means that it may be unclear whether or how the terms of existing OSS licenses apply to scenarios in which OSS projects are used to train, develop, and distribute AI systems.
@@
In the past, updated versions of popular licenses have been drafted to account for new technology (Team, 2021; Stallman, [n. d.]), and AI model-specific licenses (e.g., OpenRAIL (Responsible AI, 2022)) have emerged that impose ethical usage requirements. Thus, the new legal challenges introduced by large-scale training and GenAI may necessitate the drafting of new OSS licenses or the revision of existing licenses so as to provide guidance on these issues and allow developers to more clearly indicate their views on the use of their work in connection with GenAI. Importantly, since most existing OSS licenses were drafted by developers, who typically do not have legal experience, their terms are sometimes difficult to align with legal doctrine, giving rise to a perceived need for drafting organizations to issue FAQs explaining the intent of their licenses (Wintersgill et al., 2024). If those with legal expertise, preferably with accompanying SE backgrounds, participate in drafting these new licenses, it is possible that ambiguous and unclear language can be avoided to the largest extent possible.
8.2. Disclosing training data
Participants who had a view about the use of their code in training data communicated that it was important for licensing terms to be honored, including attribution requirements (Findings 5.2.2, 5.2.2, and 5.2.3). Absent further recognition of attribution rights under U.S. copyright law, licensing continues to be a key method for creators to enforce this preference. For this to occur, however, the content and provenance of training data must be made available to users of the models, including information about the source of each piece of training data, the license or licenses under which the data was originally issued, and the historical record of the data throughout various processes.
While regulators should consider the importance of attribution to creators, they should also recognize that any legal requirement to provide complete information on the content in training material will involve significant technical hurdles. DataBOMs (Data Bills of Materials) (Stalnaker et al., 2024; Barclay et al., 2019) are a conceptual solution but cannot easily be deployed at the scale required to capture the billions, if not trillions, of data points on which LLMs are trained  (Lee et al., 2023b). Additionally, any such provenance information must take into account that a work might appear in different ways: on a site created by the developer, in a StackOverflow post authored by another user, or as part of another developer’s work. A single DataBOM is not likely to capture these various uses. Indeed, as explained in Section 2.1, the EU AI act calls for the transparency of models and, therefore, recommends that training datasets be disclosed.
@@ 8.2. Disclosing training data
Finally, some respondents favored the recognition of a right to be removed from training data (as foreseen by legislation such as the EU AI Act). Apart from the difficulty of determining whether a work is included in a training dataset, the further question of what it means to be “removed” and how such removal can be effectively accomplished remain unanswered (and perhaps unanswerable, absent wholesale destruction of the training data)  (Lee et al., 2023a; Yao et al., 2023). This reality may align with some respondents’ views that the use of this technology may outpace at least some attempts to regulate it.
8.3. Tracking data provenance
As described above, tracking GenAI’s provenance of data output is technologically challenging. While such functionality has been achieved by newer models, like ChatGPT-4 and Gemini, this relies on Retrieval Augmented Generation (RAG)-style systems (Ma et al., 2024), which essentially depend on a database or web search to provide additional context to the model. Models employing RAG extract and/or summarize content obtained from these searches and then provide provenance information for the search results used to generate the final output. This is distinct from providing provenance information for data points that were included in the training data.
The process of retrieving provenance for these model generations that do not rely on external data sources accessible through RAG or provided in a prompt is still an open research area. As such, absent external, supplied provenance information, models may not be able to cite a specific source from which their output was derived. However, such information could be useful to address several concerns raised by developers, including being notified of information leakage (Finding 6.3) or determining the root of bugs or vulnerabilities (Finding 6.1).
Since new LLM models can be created by fine-tuning existing base models, it also becomes necessary to track the provenance and lineage of models appropriately. Failure to do so could result in not all training data being known or disclosed. These relationships emerging from the AI supply chain could be tracked using standardized model cards (Mitchell et al., 2019) or AIBOMs (AI Bills of Materials) (Xia et al., 2023; Stalnaker et al., 2024), although these methods add their own compliance challenges.
Tracking the provenance of generated code naturally requires knowledge of the generated code’s presence. However, developers indicated that the use of GenAI to produce code is rarely documented (Finding 4.2.2), and even in cases where such documentation exists, it often only notes that GenAI was utilized without providing details of the prompt or the rest of the interaction (Finding 4.2.2). This could obscure the presence of generated code, complicating provenance and making it more difficult to achieve developers’ goals surrounding attribution, license compliance, and risk mitigation concerning generated code. Future work investigating the process by which GenAI is utilized in software engineering may help to identify methods by which to facilitate documentation of GenAI usage that are useful for provenance efforts.
@@ 8.4. New AI advancements will likely amplify legal challenges
Many of the study participants expressed the sentiment that “tools are tools” and, as such, developers should hold the copyright to GenAI output (Finding 5.4.2). Much of this understanding is based on the current limitations of existing tools, which, in many cases, require developers to manually correct, edit, or adapt the solution provided by the AI before the code can be used. A few respondents, however, indicated that they anticipated more complicated legal challenges ahead once the technology advances and can produce entire software packages composed of multiple files that largely resemble existing licensed components, a process already under development (Wu, 2024; gpt-pilot, [n. d.]). The question will then be whether developers’ attitudes about generated code ownership will change based on the nature and scope of developers’ own contributions. Likewise, while developers’ intuitions about code usage may currently align, at least to some extent, with U.S. copyright law, there may well be a future misalignment as both the technology and the law evolve. Finally, our study indicates that incentives relating to creativity and the use of preexisting work might vary considerably depending on the community’s norms, such as whether the work is open-source or proprietary. Having a better understanding of developer expectations and perceptions could help guide regulators and decision-makers in this area and call into question whether “one size fits all” is the proper approach.
Several study participants also emphasized the speed at which AI development is taking place and whether law and regulation would be able to keep up (Finding 5.1). Other respondents conveyed the view that GenAI was different only in degree and not in kind from other creative processes that did not attract regulation. Still others reported confidence in their views on as-yet unresolved legal issues regarding copyright and GenAI. Such respondents, collectively, seemed to suggest that the law may not be a strong motivator in this development space. Decision makers might, therefore, consider whether and how regulation can address concerns without being seen as an obstacle merely to be designed around.
Finally, to the extent that copyright law is concerned with incentivizing creativity, GenAI regulators should not think of copyright law in a vacuum. The other legal concerns raised by developers—such as concerns about privacy, security, and tort liability (Finding 6.1)—are likely to have an effect on how and when they use GenAI and may, in fact, work at cross-purposes with the motivations underlying copyright law. (For example, developers might be conflicted about asserting authorship over GenAI work for copyright purposes if that means also bearing responsibility for unforeseen harms.)
8.5. Final thoughts and implications
@@
Although there are many open questions regarding the use of generative AI and copyright, the biggest priority for future research is addressing problems related to data provenance. Developers expect tools that can provide them with citations and will not lead them unwittingly into legal trouble. Likewise, they need tools that allow them to document and track code generated by LLMs for transparency and control of legal risk. Given that this paper primarily examines these issues from a perspective grounded in U.S. law, we encourage further research to interpret our results against regulatory efforts elsewhere.
9. Conclusion
In this paper, we presented a study conducted by a joint team of SE and legal researchers, which surveyed 574 software developers from across the globe who use GenAI tools for various software development tasks. Through an online survey and follow-up interviews, we explored developers’ perspectives on emerging legal issues, including their perception of copyrightability, ownership of generated code, and other related concerns. Our aim was to identify potential misconceptions among developers, assess the impact of GenAI on their work, and evaluate their awareness of licensing and copyright risks. By using both qualitative and quantitative methods to analyze the responses, we found that opinions on copyright issues vary widely, especially regarding the ownership of model outputs. Additionally, many developers demonstrated an awareness of the complexities and nuances involved in addressing these increasingly complex legal questions. Finally, we provided discussion on our findings that can inform policy decisions and guide future research.
10. Acknowledgments
We would like to thank the study participants for their time and valuable contributions. This research has been supported in part by NSF grant CCF-2217733. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors. A complete, detailed list of image attributions can be found in our online replication package (ano, 2024).